Tavon has a gift card for $85 that loses $2 for each​ 30-day period it is not used. He has another gift card for $40 that loses $1.50 for each​ 30-day period it is not used.
1
Expert's answer
2020-09-21T14:04:15-0400



Let k be the number of 30day period

85-2k = 40-1.50k

85-40 = -1.50k +2k

45 = 0.50k

k= 90.

Tavon will prefer the cards at 90th 30-day period.